94)
Sharing cost. The members of a flying club plan to share
equally the cost of a $200,000 airplane. The members
want to find five more people to join the club so that the
cost per person will decrease by $2000. How many members
are currently in the club?

Answer by stanbon(75887)

Sharing cost. The members of a flying club plan to share
equally the cost of a $200,000 airplane. The members
want to find five more people to join the club so that the
cost per person will decrease by $2000. How many members
are currently in the club?
-----------
Let original number of members be "x".
Original average share cost: 200,000/x
---------------
New number of members: "x+5"
New average share cost: 200,000/(x+5)
------------------
Equation:
original share cost - new share cost = 2000
200,000/x - 200,000/(x+5) = 2000
100/x - 100/(x+5) = 1
100(x+5) - 100x = x(x+5)
500 = x^2 + 5x
x^2 + 5x - 500 = 0
(x+25)(x-20) = 0
Positive answer:
x = 20 (original number of members)
===================================
Cheers,
Stan H.